# On The Heine-Borel Property and Minimum Enclosing Balls

### v1

@@ 4 Hilbert Radius Minimum Enclosing Ball
We begin by showing that the Hilbert metric spaces have the minimum ball property.
Lemma 4.1.
Any Hilbert metric space $(\Omega,d)$ has the Heine-Borel property.
Proof 4.2.
This follows from Theorem 2.1 in [3] where they show that the Hilbert and Euclidean metrics induce the same topology on bounded convex domains in $\mathbb{R}^{n}$.
Since every Hilbert metric space satisfies the minimum ball property, we know by Theorem 3.7 the minimum radius enclosing ball problem is LP type in a Hilbert metric. By Proposition 3.10 the well order $\preceq$ on $\mathbf{X}$ can be replaced with the lexicographical usual total order $\leq$ from $\mathbb{R}^{k}$. This is useful as we can comprehend and compute using the lexicographic total order.
Running an LP algorithm on an LP-type problem requires the solving of two primitive operations [16]. These are the violation test (given a basis $B\subset S$ and element $x\in S$, whether $f(B)=f(B\cup\{x\})$) and basis computations (how to find a basis of $B\cup\{x\}$). Instead of a well-order, we will let $\delta$ be the lexicographic order $\mathbb{R}^{2}$ for the Hilbert convex polygons. We need the two following supporting lemmas:
@@
This follows immediately from the fact that, in general position, at most three points defines a Hilbert ball [10].
Lemma 4.5.
Given an $m$-sided convex polygon $\Omega$ in $\mathbb{R}^{2}$, the center of the minimum Hilbert radius ball around two points, $p,q\in\operatorname{int}\Omega$, can be computed in time $O(\log m)$.
Proof 4.6.
The balls of minimum radius for a set of two points are always centered on a section of the bisector in the sector directly between the two points. This is because Hilbert balls are polygons and the balls around two points meet at a segment across this sector [18]. This piece of the bisector is a line [4]. We perform binary search the boundary of $\Omega$ to determine which edges define this sector, and then use the bisector equation there [9, 4]. We choose the lexicographically smallest point along the bisector in this sector to serve as the center of the minimum Hilbert radius ball around points $p,q$.
@@
Case 1: $B$ has two elements $y,z$. To determine if $f(B)=f(B\cup\{x\})$, it suffices to check whether $x$ is contained in the minimum radius ball of $y,z$. If $x$ is contained, then $f(B)=f(B\cup\{x\})$. If $x$ is not contained, either the center must move or the radius must increase, so $f(B)<f(B\cup\{x\})$. Since we can compute the center of a two-point Hilbert ball in $O(\log m)$ time by Lemma 4.5, we can check this case in $O(\log m)$ time.
Case 2: $B$ has three elements. In this case, we calculate the center of the Hilbert ball around the three points in time $O(\log^{3}m)$ using the algorithm given in [9]. If a center exists, we compute the distance between $x$ and the center. If a center does not exist, one of the three elements is contained in a ball defined by the other two, so we check all pairs of elements using Lemma 4.5 in $O(\log m)$ time to find the minimum enclosing ball. We take the resulting ball and check the center’s distance to $x$. Either $x$ is in the ball or it is not, in which case $f(B)$ must increase because either the radius or center must move.
@@
Suppose we have a previous basis $B$ and a new element $x$ that is not contained inside the ball formed by the basis. We would like to compute the basis of $B\cup\{x\}$. This gives us a few cases. We will consider the non-trivial cases where $B$ contains two or three elements. In either situation, we compute the minimum enclosing balls of all pairs and triples of points in $B\cup\{x\}$, and check for containment of the remaining points in those balls. This can be done in $O(\log^{3}m)$ as described in Case 2 in the proof of Lemma 4.7.
Theorem 4.11.
Hilbert minimum radius balls can be computed in time $O(n\log^{3}m)$.
Proof 4.12.
This follows directly from the running time of our primitve operations and Theorem 7 and Corollary 8 from Sharir and Welzl’s “A combinatorial bound for linear programming and related problems” [25].
5 Thompson and Funk Minimum Radius Enclosing Ball
Given Theorem 3.7 we can show that minimum radius balls in the Thompson metric are also LP type. This proof follows the format of Proposition 6.1 in [21].
Figure 3: A Thompson ball of radius 1 (outlined in black), nested between two Hilbert balls of radius 1 and $1/2$ (outlined in red), all centered at $p$.
Theorem 5.1.
The topology induced by the Thompson metric in a bounded convex domain $\Omega\subset\mathbb{R}^{d}$ coincides with the Hilbert topology in that domain.
Proof 5.2.
To do this, we will compare balls in both geometries. Given a convex body $\Omega\subset\mathbb{R}^{d}$, for every point $p\in\Omega$ and radius $0<r$, we will show:
$B_{H}(p,r/2)\leavevmode\nobreak\ \subset\leavevmode\nobreak\ B_{T}(p,r)\leavevmode\nobreak\ \subset\leavevmode\nobreak\ B_{H}(p,r).$
For an example of this, see Figure 3. To show that $B_{T}(p,r)\subset B_{H}(p,r)$, we must show that the Thompson distance from $p$ to a point on the boundary of the ball is greater than or equal to the Hilbert. This follows directly from the definition of the Hilbert and Thompson metrics:
$H(a,b)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{2}(F(a,b)+F(b,a))\leq 2\cdot\frac{1}{2}(\max(F(a,b),F(b,a))=T(a,b).$
To prove that $B_{H}(p,r/2)\subset B_{T}(p,r)$, we must show that the Thompson distance from $p$ to a point on the boundary of the ball is less than or equal to the Hilbert distance. This follows from the fact that if $H(a,b)=\frac{1}{2}r$ then $\frac{1}{2}(F(a,b)+F(b,a))=\frac{1}{2}r$ so $F(a,b)+F(b,a)=r$ and so $T(a,b)=\max(F(a,b),F(b,a))\leq r$.
Hence, the family of balls $B_{H}(p,r)$ and $B_{T}(p,r)$ are subbases for the same topology.
Corollary 5.3.
The minimum radius ball problem in the Thompson metric is LP type.
Proof 5.4.
Since both metric spaces have the same topology, and we know that the Hilbert metric satisfies the Heine-Borel property, we know that the Thompson must as well. Hence, the Thompson metric has the minimum ball property, so computing minimum radius balls in the Thompson metric are LP type.
Proposition 5.5.
Minimum radius balls in the Funk, reverse Funk weak metrics are LP type.
Proof 5.6.
This follows from two facts, the first being that Proposition 6.1 in [21] proves that in a bounded convex $\Omega$ the funk and reverse funk metrics have the same topology as the Euclidean topology in that domain, thus proving it has the Heine-Borel property[21]. Secondly, as shown in Corollary 3.5 and Corollary 3.9 this means that minimum radius balls are LP type in the Funk metrics.
6 Conclusions
In this paper, we contributed a criterion for showing that minimum radius balls are LP type. This criterion being that finite sets in the metric space always have at least one minimum radius ball. We showed that if a metric space, or weak metric space, has the Heine-Borel property, it has this property. We used this to contribute a minimum radius ball algorithm for the Hilbert metric and proved that the minimum radius ball problem is LP type for the Thompson metric and Funk weak metrics. Additionally, we proved that the topology induced by the Thompson metric coincides with the Hilbert, and that balls in the two metric are equivalent up to scaling factors.

### v2

@@ 4 Hilbert Radius Minimum Enclosing Ball
We have shown in Corollary 3.12 that the minimum radius enclosing ball problem is LP-type in a Hilbert metric, so we will now focus on how to implement the LP algorithm. By Proposition 3.10 the well-order $\preceq$ on $\mathbf{X}$ can be replaced with the lexicographical usual total order $\leq$ from $\mathbb{R}^{k}$. This is useful as we can comprehend and compute using the lexicographic total order.
Running an LP algorithm on an LP-type problem requires solving two primitive operations [17]. These are the violation test (given a basis $B\subset S$ and element $x\in S$, whether $f(B)=f(B\cup\{x\})$) and basis computations (how to find a basis of $B\cup\{x\}$). Instead of a well-order, we will let $\delta$ be the lexicographic order $\mathbb{R}^{2}$ for the Hilbert convex polygons. We need the two following supporting lemmas:
@@
Proof 4.2.
This follows immediately from the fact that, in general position at most three points defines a Hilbert ball [9](see Lemma 14). This is because Hilbert balls intersect along line segment edges in general position. Three points cannot intersect along the same edge at the same distance without two of them lying along the same spoke.
Lemma 4.3.
Given an $m$ sided convex polygon $\Omega$ in $\mathbb{R}^{2}$, the center of the minimum Hilbert radius ball around two points, $p,q\in\operatorname{int}\Omega$, can be computed in time $O(\log m)$.
@@
The balls of minimum radius for a set of two points are always centered on a section of the bisector in the sector directly between the two points. This is because Hilbert balls are polygons and the balls around two points meet at a segment across this sector[19] (see Lemma $10$ page $4$). This piece of the bisector is a line [4]. We binary search the boundary of $\Omega$ to determine which edges define this sector, and then use the bisector equation there [9, 4] (see Section 3 in [9]). We choose the lexicographically smallest point along the bisector in this sector to serve as the center of the minimum Hilbert radius ball around points $p,q$.
@@
Case 1: $B$ has two elements $y,z$. To determine if $f(B)=f(B\cup\{x\})$, it suffices to check whether $x$ is contained in the minimum radius ball of $y,z$. If $x$ is contained, then $f(B)=f(B\cup\{x\})$. If $x$ is not contained, either the center must move or the radius must increase, so $f(B)<f(B\cup\{x\})$. Since we can compute the center of a $2$ point Hilbert ball in $O(\log m)$ time by Lemma 4.3, we can check this case in $O(\log m)$ time.
Case 2: $B$ has three elements. In this case, we calculate the center of the Hilbert ball around the three points in time $O(\log^{3}m)$ using the algorithm from "Delaunay Triangulations in the Hilbert Metric" [9]. If a center exists, we compute the distance between $x$ and the center. If a center does not exist, one of the three elements is contained in a ball defined by the other two, so we check all pairs of elements using Lemma 4.3 in $O(\log m)$ time to find the minimum enclosing ball. We take the resulting ball and check the center’s distance to $x$. Either $x$ is in the ball or it is not, in which case $f(B)$ must increase because either the radius or center must move.
@@
Proof 4.8.
Suppose we have a previous basis $B$ and a new element $x$ that is not contained inside the ball formed by the basis. We would like to compute the basis of $B\cup\{x\}$. This gives us a few cases. We will consider the non-trivial cases where $B$ contains two or three elements. In either situation, we compute the minimum enclosing balls of all pairs and triples of points in $B\cup\{x\}$, and check for containment of the remaining points in those balls. This can be done in $O(\log^{3}m)$ as described in Case 3 of Lemma 4.5.
Theorem 4.9.
Hilbert minimum radius balls, of $n$ points, can be computed in time $O(n\log^{3}m)$.
@@
This follows directly from the running time of our primitive operations and Theorem 7 and Corollary 8 from Sharir and Welzl’s "A combinatorial bound for linear programming and related problems"[26].
5 Conclusion
In this paper, we contributed a criterion for showing that minimum radius balls are LP-type. This criterion being that finite sets in the metric space always have at least one minimum radius ball. We showed that if a metric space, or weak metric space, has the Heine-Borel property, it has this property. We used this to contribute a minimum radius ball algorithm for the Hilbert metric and proved that the minimum radius ball problem is LP-type for the Thompson metric and Funk weak metrics.